A 60.2 kg skier rides a 831.5 m long lift to the top of a mountain. The lift makes an angle of 16.8o with the horizontal. How much work needs to be done on the skier to get her to the top of the mountain?
1
Expert's answer
2020-02-10T09:07:02-0500

Given that mass =60.2kg

Length of the lift=831.5 m

And, lift is making "\theta=16.8^\circ"

So, work done ="mgh"

"W=60.2 \times 9.8 \times 831.5 \sin 16.8 =141785.05J"